Week 14 of the Early Operations system optimization period
Following stabilization of the dynalene coolant flow to the telescope top end assembly, and restoring the LSST Camera to an operational vacuum level and focal plane temperature, on-sky testing resumed on the night of 27 January. Together with the LSST Camera restart, the team deployed the previously tested sequencer CCD readout controller firmware update to adopt the natural frequency of the readout electronics for regular nighttime operations.
The first night back on sky was largely devoted to two engineering tests to evaluate the delivered image quality around filter changes. First, re-measuring optimal filter focus offsets by cycling between the r and y filters, and second, a test to isolate the effects of filter changes and rotator angle movements (for context, filter changes require the rotator to move to the zero degree position for vertically moving the filters during the exchange). An updated filter focus offset for the y filter has now been deployed.
The team used pre-LSST survey observations to validate recent development efforts to speed up the production of active optics system (AOS) closed loop corrections. The recent work reduced the time between the shuttler close and the start of image reduction for the wavefront sensors. The current typical time between shutter close and delivery of the wavefront corrections is ~40 seconds. Rubin Observatory is approaching the capability to routinely apply optical corrections determined from visit N to visit N+2, which requires the corrections to be computed within ~35 seconds from shutter close.
The team also deployed an update for the AOS closed loop control system to specify the gain for each individual basis function in the space of basis functions that minimizes degeneracies between degrees of freedom. This update is intended to enhance simultaneous control of de-focus and spherical optical aberrations in the near term, and provide increased flexibility for further optimization.
A filter swap to remove the u filter and install the y filter occurred on 27 January. The currently available filter set during lunar bright time is grizy.